OpenStudy (anonymous):

A drawer contains 3 tank tops, 6 short-sleeve shirts, 4 three-quarter sleeve shirts, and 5 long-sleeve shirts. What is the probability of randomly choosing a tank top and then a three-quarter sleeve length, if you replace the tank top?

OpenStudy (anonymous):

3+6+4+5=18 items

OpenStudy (anonymous):

8/18

OpenStudy (anonymous):

my answers dont go that high

OpenStudy (anonymous):

7/18, 1/27, 2/3, or 1/18

OpenStudy (anonymous):

Sorry \[ \frac 3 {18} \frac4{18}=\frac 1{27} \]

OpenStudy (anonymous):

oh alright. ty

OpenStudy (anonymous):

yw. There are 4 three-quarter sleeve shirts